# Federated Learning for Traffic Flow Prediction with Synthetic Data Augmentation

### v1

@@ 4.2. Federated Traffic Flow Prediction Model
(7) $f_{c}(\omega)=\frac{1}{|Q_{c}|}\sum_{t}(\mathcal{L}(\mathcal{F}_{c}(\omega;q_{t-h:t}),q_{t+\tau})$
@@
(8) $\textnormal{argmin}f(\omega):=\sum^{C}_{c=1}\frac{|\bm{Q_{c}}|}{|Q_{\textnormal{glob}}|}f_{c}(\omega)$
@@
11:         Global updates $\omega_{r_{\textnormal{glob}}+1}\leftarrow\sum_{c=1}^{C}\frac{|\bm{Q_{c}}|}{|Q_{\textnormal{glob}}|}f(\omega)_{r_{\textnormal{glob}}}^{c}$
@@
4.3. Proposed Traffic Flow Prediction Model
The Temporal Attention Unit was introduced by Tan et al (Tan et al., 2023) and poses temporal attention as a combination of statical attention, which occurs within a data sample, and dynamical attention, which occurs between the frames of the data samples. The statical attention can be captured by performing depth-wise convolution, followed by a dilated depth-wise convolution, and finally a 2D convolution. The depth-wise convolution layers learn separate kernel filters for each of the channels in the encoded representation of the input. The dynamical attention is composed of a 2D average pooling layer, followed by a fully-connected layer. The final temporal attention is a result of the product between the statical and dynamical attention components.
An Encoder-Decoder architecture model architecture is presented in this work, wherein the Encoder and Decoder represent stacked 2D convolutional layers that encode a spatial representation of the data. To augment this Encoder-Decoder architecture for the task of traffic prediction, we can model the regional traffic flow data as a sequence of graphs, and employ the Graph Attention model (Veličković et al., 2018) with multi-head attention, to learn multi-channel relationships between regions in the city. Figure 4 compares the model architecture between TAU and GATAU.
Figure 4. General model architecture comparison of TAU and GATAU.
[…]
Model architecture comparison
Thus, we introduce the Graph Attention with Temporal Attention Unit (GATAU). Figure 5 illustrates the GATAU module in further detail.
Figure 5. Architecture of the proposed GATAU model.
[…]
Architecture of our spatio-temporal model
Our input data $X\in\mathbb{R}^{N\times T}$, where $N$ is the number of nodes and $T$ is the length of the input sequence, is first processed by a 2D convolutional layer to map the input into a higher dimension feature space. This output is fed into the Multi-Head Graph Attention (MGAT) module. Within the MGAT, attention coefficients are calculated which give a measure of the importance of relationships between nodes. The input into MGAT are node features, $h=\{\vec{h}_{1},\vec{h}_{2},...,\vec{h}_{N}\}\in\mathbb{R}^{T\times F}$ where $F$ is the input feature size. Attention coefficients for layer $l$ are calculated by
(9) $e_{ij}^{l}=\textnormal{LeakyReLU}\left(\vec{a}^{l^{T}}\left[W^{l}\vec{h}^{l}_{i}||W^{l}\vec{h}^{l}_{j}\right]\right)$
@@ 4.3. Proposed Traffic Flow Prediction Model
(10) $a_{ij}^{l}=\frac{\textnormal{exp}(e^{l}_{ij})}{\sum_{k\in\mathcal{N}_{i}}\textnormal{exp}(e_{ik}^{l})}$
@@
(11) $h_{i}^{l+1}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}W^{k}\vec{h}_{j}\right)$
where $\sigma$ is an ELU activation function. To enhance the stability of the model during training, a multi-head mechanism of attention is utilised for the graph attention module, wherein each attention head trains its own parameters. As per the original work which averages the multi-head outputs at the final layer (Veličković et al., 2018), our model also averages the multi-head outputs after the single-layer MGAT module. The output from the MGAT module is of shape $X\in\mathbb{R}^{N\times T\times F^{\prime}}$.
A residual connection is used to fuse the output of the initial convolution layer with the output from the MGAT module. The fused hidden layer is then reshaped into $X\in\mathbb{R}^{N\times F^{\prime}\times H\times W}$ format to be processed by the TAU module, where $H$ and $W$ are the dimensions of the rectangular grid of the city. Our output layer is composed of a convolutional layer, and fully connected layer, which reduce the dimensionality of our processed input’s feature space to match our target shape, which is $Y\in\mathbb{R}^{N}$.
5. Experiment
5.1. Dataset
The experiments conducted in our work use the ride-sharing dataset published by Didi Chuxing GAIA Initiative. The dataset contains data for over three million journeys made in the city of Chengdu, China over the entire month of November 2016. Table 1 contains details of the dataset.
Table 1. Description of Didi Chuxing dataset for the city of Chengdu.
Dataset Chengdu
Orders 5,807,001
Drivers 1,136,970
Latitudes (30.65293, 30.72776)
Longitudes (104.042135, 104.12959)
@@ 5.2. Data Pre-Processing
Targeting the real-world setting of multiple organisations collaborating to train trajectory generation and traffic flow prediction models through FL, the Didi Chuxing dataset is partitioned into subsets based on the number of clients emulated in the FL setting. It is assumed that each vehicle will collect data for a particular organisation, and hence, trajectories will be grouped according to the driver vehicle, and vehicles will be assigned exclusively to a particular organisation partition. For each client setting we partition vehicles evenly among the clients. Since individual drivers in a ride-sharing context will have preferences over their journeys and the number of orders they are assigned this will embed feature-based skews across the partitions and ensure realistic data distributions.
Once partitioned, we employ the same pre-processing steps as (Zhu et al., 2023) to train the trajectory generation model. Each client also processes their trajectory dataset into regional vehicle inflow format for the traffic flow prediction task. For this step, the city of Chengdu is discretised into a rectangular 10x10 grid. Regional inflows are aggregated into 30 minute windows, and sequences of these windows over 6 hours are joined to be used by the traffic prediction model to predict the regional inflow 3 hours in the future. For the traffic flow prediction task, we segment the regional inflow data into the training, evaluation, and testing datasets by splitting the entire dataset into sets of 60%, 20%, and 20%, respectively.
@@
As a measure of the heterogeneity between the regional inflow datasets, we employ two image-based comparison metrics to assess average similarity across the partitioned client datasets: normalised Root Mean Squared Difference (nRMSD) and Structural Similarity Index (SSI) - whose value ranges from 1 (identical) to -1 (dissimilar). The properties of the partitions detailed in Table 2 suggest that as the number of clients in a setting increases, data heterogeneity also increases; both metrics show a decrease in data similarity as the client number increases.
@@
Table 2. Metrics assessing the average similarity between the partitioned datasets representing each client’s local dataset for each FL setting.
Clients # Orders per Client Partition nRMSD SSI
2 (2904846, 2902155) 0.15 0.95
4 (1452404, 1452464, 1453202, 1448931) 0.21 0.92
6 (969168, 965928, 969817, 968308, 969601, 964179) 0.26 0.88
5.3. Description of Experiments
  1. (1)
@@ 5.3. Description of Experiments
    Trajectory Data Generation: This experiment details the performance of the trajectory generation models as a function of data heterogeneity and number of clients. This is important as it demonstrates the generative capability of a trajectory generation model in federated implementations with greater data heterogeneity..
  2. (2)
@@
  3. (3)
    Pre-Training with Synthetic Data: This study explores the effects of pre-training the global model in an FL framework using synthetic data. This is important, as it provides a method for training traffic flow prediction models in scenarios with limited communication or local training rounds.
  4. (4)
    Single Client Data Augmentation: In this experiment, the single client case is studied, wherein all data is owned by one organisation. The synthetic data generation method from FedTPS can be employed by the client to augment their dataset. This study emphasises the benefits of increasing the dataset size on the training process of a client.
  5. (5)
@@
5.4. Baselines
Unless otherwise stated, all models are trained using the Adam optimiser with a learning rate of 0.001 and a batch size of 32. For federated training, we use 80 global training rounds where we aggregate the updated client parameters after one local training round. The following are the settings of the various baselines for the traffic flow prediction task, including our proposed model GATAU.
  1. (1)
    GRU: A three-layer GRU model (Cho et al., 2014) with hidden dimensions of 32 is implemented.
  2. (2)
    STGCN: The STGCN model (Yu et al., 2018) is implemented with the same model parameters specified in the original work. A distance based adjacency matrix is used for the Graph Convolution operation.
  3. (3)
    DCRNN: The DCRNN model (Li et al., 2018) is implemented with the same model parameters specified in the original work. A distance based adjacency matrix is used for the Diffusion Convolution operation.
  4. (4)
    GWNET: The GWNET model (Wu et al., 2019) is implemented with the same model parameters specified in the original work. For the self-adaptive adjacency matrix, the input graph assumes connections for regions within 4 kilometres.
  5. (5)
    TAU: The TAU model (Tan et al., 2023) is implemented with the same model parameters specified in the original work.
  6. (6)
    GATAU: The input convolutional layer is composed of two sequential 2D convolution layers with a hidden size 16. The TAU units have the same model parameters as in the original work (Tan et al., 2023). Our output layer is composed of two 2D convolution layers at output with hidden size 16, and a kernel size (1,1), one final linear layer to match dimensions to the output dimension. For the Graph Attention layer, the graph assumes connections between regions within 4 kilometres.
@@ 5.4. Baselines
  2. (2)
    FedAvg: FedAvg (McMahan et al., 2017) uses a weighted averaging mechanism, based on relative client dataset sizes, to update the global model based on individual client contributions.
  3. (3)
    FedOpt: FedOpt (Reddi et al., 2021) builds on FedAvg by implementing server-side adaptive optimisers such as Adam, as well as server-side momentum. For the server, we use a learning rate of 0.01, a momentum parameter of 0.9, and a second moment parameter of 0.99.
  4. (4)
    FedProx: FedProx (Li et al., 2020) utilises a penalty term, $\mu$, to inhibit deviations of clients from the global model. We use a proximal term $\mu$ of 0.001.
  5. (5)
@@
A server with two NVIDIA RTX 3090 GPUs is used to conduct our experiments. Our code, as well as a list of the pre-requisite packages, is made available at https://anonymous.4open.science/r/FedTPS-C51B/.
6. Results
6.1. Trajectory Data Generation
Figure 6 depicts trajectories generated by the trajectory generation model for a variety of different client settings. The single client example is equivalent to a centralised training paradigm. As the number of clients increases, the generated trajectories can be observed to be noisier. After converting the generated trajectories into regional inflow data, table 3 shows the similarity between generated samples and real traffic data based on on three metrics, averaged over the regional inflow data sequences. Some deviation is expected due to the generational diversity of the DiffTraj model, yet the results show close similarity meaning that the spatio-temporal characteristics of the original data were modelled accurately by the trajectory generation model. In general, increasing the number of clients in the FL setting results in a larger deviation from the real traffic data for the samples generated, as evidenced by the decreasing values of SSI and increasing MAD and nRMSE metrics.
Table 3. Comparison of generated regional inflow data against real regional inflow data for the DiffTraj model trained through FL, with varying client cases.
City Clients MAD nRMSE SSI
Chengdu 1 11.03 0.10 0.98
2 12.24 0.10 0.98
4 12.91 0.11 0.97
6 13.30 0.11 0.96
Table 4. Comparison of various Spatio-Temporal model performances for prediction of traffic at 6 time-steps in future. Assessed using nMAE and MAPE metrics, under varying client numbers and varying FL methods.

### v2

@@ IV-B Federated Traffic Flow Prediction Task
$f_{c}(\omega)=\frac{1}{|Q_{c}|}\sum_{t}(\mathcal{L}(\mathcal{F}_{c}(\omega;q_{t-h:t}),q_{t+\tau})$ (7)
@@
$\textnormal{argmin}f(\omega):=\sum^{C}_{c=1}\frac{|Q_{c}|}{|Q_{\textnormal{glob}}|}f_{c}(\omega)$ (8)
@@
11:         Global updates $\omega_{r_{\textnormal{glob}}+1}\leftarrow\sum_{c=1}^{C}\frac{|Q_{c}|}{|Q_{\textnormal{glob}}|}f(\omega)_{r_{\textnormal{glob}}}^{c}$
@@
IV-C Proposed Traffic Flow Prediction Model
The Temporal Attention Unit was introduced by Tan et al [5] and poses temporal attention as a combination of statical attention, which occurs within a data sample, and dynamical attention, which occurs between the frames of the data samples. The statical attention can be captured by performing depth-wise convolution, followed by a dilated depth-wise convolution, and finally a 2D convolution. The depth-wise convolution layers learn separate kernel filters for each of the channels in the encoded representation of the input. The dynamical attention is composed of a 2D average pooling layer, followed by a fully-connected layer. The final temporal attention is a result of the product between the statical and dynamical attention components.
An Encoder-Decoder architecture model architecture is presented in this work, wherein the Encoder and Decoder represent stacked 2D convolutional layers that encode a spatial representation of the data. To augment this Encoder-Decoder architecture for the task of traffic prediction, we can model the regional traffic flow data as a sequence of graphs, and employ the Graph Attention model [28] with multi-head attention, to learn multi-channel relationships between regions in the city. Figure 3 compares the model architecture between TAU and GATAU.
Figure 3: General model architecture comparison of TAU and GATAU.
Thus, we propose the Graph Attention with Temporal Attention Unit (GATAU). Figure 4 illustrates the GATAU module in further detail.
Figure 4: Architecture of the proposed GATAU model.
Our input data $Q\in\mathbb{R}^{N\times T}$, where $N$ is the number of nodes and $T$ is the length of the input sequence, is first processed by a 2D convolutional layer to map the input into a higher dimension feature space. This output is fed into the Multi-Head Graph Attention (MGAT) module. Within the MGAT, attention coefficients are calculated which give a measure of the importance of relationships between nodes. The input into MGAT are node features, $h=\{\vec{h}_{1},\vec{h}_{2},...,\vec{h}_{N}\}\in\mathbb{R}^{T\times F}$ where $F$ is the input feature size. Attention coefficients for layer $l$ are calculated by
$e_{ij}^{l}=\textnormal{LeakyReLU}\left(\vec{a}^{l^{T}}\left[W^{l}\vec{h}^{l}_{i}||W^{l}\vec{h}^{l}_{j}\right]\right)$ (9)
@@ IV-C Proposed Traffic Flow Prediction Model
$a_{ij}^{l}=\frac{\textnormal{exp}(e^{l}_{ij})}{\sum_{k\in\mathcal{N}_{i}}\textnormal{exp}(e_{ik}^{l})}$ (10)
@@
$h_{i}^{l+1}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}W^{k}\vec{h}_{j}\right)$ (11)
where $\sigma$ is an ELU activation function. To enhance the stability of the model during training, a multi-head mechanism of attention is utilised for the graph attention module, wherein each attention head trains its own parameters. As per the original work which averages the multi-head outputs at the final layer [28], our model also averages the multi-head outputs after the single-layer MGAT module. The output from the MGAT module is of shape $[N,T,F^{\prime}]$.
A residual connection is used to fuse the output of the initial convolution layer with the output from the MGAT module. The fused hidden layer is then reshaped into $[N,F^{\prime},H,W]$ format to be processed by the TAU module, where $H$ and $W$ are the dimensions of the rectangular grid of the city. Our output layer is composed of a convolutional layer, and fully connected layer, which reduce the dimensionality of our processed input’s feature space to match our target shape, which is $Y\in\mathbb{R}^{N}$.
V Experiments
V-A Dataset
The experiments in this study utilise the ride-sharing dataset published by Didi Chuxing GAIA Initiative, spanning the cities of Chengdu and Xi’an in China during November 2016. Chengdu has a population of 21 million and features a ring-road-based road network, consistent traffic patterns influenced by stable weather, little prominent tourism variation, and broader spatial traffic distribution due to urban sprawl. In contrast, Xi’an has nearly 13 million residents and features a grid-based road network, greater traffic variability due to weather and tourism, and a dense historic core that concentrates traffic patterns. The different characteristics of these two datasets provide a diverse study to evaluate FedTPS’s ability to enhance traffic flow prediction models in different settings. Table I contains details of the datasets.
Figure 5: Satellite images of Chengdu (left) and Xi’an (right), with real sample trajectories shown in cyan representing $X_{c}$, and where traffic flow into the red regions represents $Q_{c}$.
TABLE I: Description of Didi Chuxing dataset for the cities of Chengdu and Xi’an.
Dataset Chengdu Xi’an
Orders 5,807,001 3,410,420
Drivers 1,136,970 520,181
Latitudes (30.65293, 30.72776) (34.2053, 34.28024)
Longitudes (104.042135, 104.12959) (108.91114, 108.9986)
@@ V-B Data Pre-Processing
The Didi Chuxing dataset is partitioned into subsets based on the number of clients emulated in the FL setting. It is assumed that each vehicle will collect data for a particular organisation, and hence, trajectories will be grouped according to the driver vehicle, and vehicles will be assigned exclusively to a particular organisation partition. For each client setting we partition vehicles evenly among the clients. Since individual drivers in a ride-sharing context will have preferences over their journeys and the number of orders they are assigned this will embed feature-based skews across the partitions and ensure realistic data distributions.
The partitioned trajectories are then processed into regional vehicle inflow format by dividing the cities into rectangular 10x10 regions, and vehicle inflows over each region are aggregated over 30 minute intervals. For the traffic flow data generation task, the partitioned traffic flow datasets, as well as the conditional dataset, are used to train TFDiff to generate synthetic traffic flow data possessing the same underlying spatio-temporal characteristics as the real data with some generational diversity. Then, FedTPS augments each client’s traffic flow data with the generated synthetic data. Sequences of traffic flow data over previous 6 hours are used by the traffic prediction models to predict the regional inflow 3 hours in the future. The regional inflow data are divided into training, evaluation, and testing datasets, with the first 75% used for training, and the final 25% for testing.
@@
As a measure of the heterogeneity between the regional inflow datasets, we employ three comparison metrics to assess average similarity across the partitioned client datasets: normalised Mean Absolute Difference (nMAD), Structural Similarity Index (SSI), and Temporal Correlation (TC) coefficient, where for each client and region, we calculate the temporal correlation between the two datasets over time. The properties of the partitions detailed in Table II suggest that as the number of clients in a setting increases, data heterogeneity also increases; both metrics show a decrease in data similarity as the client number increases.
@@
TABLE II: Metrics assessing the average similarity between the partitioned datasets representing each client’s local dataset for each FL setting.
City Clients # Orders per Client Partition nMAD SSI TC
Chengdu 2 (2904846, 2902155) 0.1017 0.966 0.932
4 (1452404, 1452464, 1453202, 1448931) 0.1434 0.936 0.884
6 (969168, 965928, 969817, 968308, 969601, 964179) 0.1754 0.908 0.846
8 (488639, 485265, 473309, 473086, 472199, 471045, 472957, 474078) 0.2022 0.883 0.812
Xi’an 2 (1635623, 1686826) 0.1283 0.916 0.936
4 (841194, 844346, 841598, 845189) 0.1808 0.851 0.888
6 (562312, 560911, 562395, 561577, 561008, 564124) 0.2215 0.799 0.846
8 (421940, 419310, 422529, 421863, 421219, 420376, 421667, 423423) 0.2553 0.757 0.809
V-C Description of Experiments
  1. 1.
@@ V-C Description of Experiments
    Traffic Flow Data Generation: These experiment detail the performance of the data generation models as a function of data heterogeneity and number of clients. This is important as it demonstrates the generative capability of the traffic flow generation models in federated implementations with greater data heterogeneity.
  2. 2.
@@
  3. 3.
    Pre-Training with Synthetic Data: This study explores the effects of pre-training the global model in an FL framework using synthetic data. This study highlights pre-training as a method for training traffic flow prediction models in scenarios with limited communication or training rounds.
  4. 4.
@@
V-D Baselines
For the traffic flow data generation task we use 300 global training rounds where we aggregate the updated client parameters after one local training round, and utilise the Adam optimiser with a learning rate of 0.0001 and a batch size of 32. We implement various baseline models including TFDiff.
  1. 1.
    CVAE: A conditional VAE with a fully connected encoder and decoder, each containing three hidden layers, and featuring a latent space dimension 20.
  2. 2.
    CGAN: A conditional GAN with three fully connected hidden layers in both the generator and discriminator, and taking a latent vector of size 100.
  3. 3.
    TFDiff: A conditional diffusion model, with a Wide & Deep network structure employed to embed conditional information. The encoder and decoder blocks in the UNet architecture feature multiple 1D-CNN-based stacked residual network blocks (Resnet block) with a kernel size of 3. An intermediate block, which is an attentive layer between two Resnet blocks with kernel size 1, is implemented between the encoder and decoder blocks.
For the traffic flow prediction task we use 80 global training rounds where we aggregate the updated client parameters after one local training round, and utilise the Adam optimiser with a learning rate of 0.001 and a batch size of 32. We use various baseline models with the same model parameters as the original work, unless otherwise stated. The baselines are implemented with the following settings:
  1. 1.
    GRU: A three-layer GRU model [6] with hidden dimensions of 32 is implemented.
  2. 2.
    STGCN: The STGCN model [2] using a distance based adjacency matrix.
  3. 3.
    DCRNN: The DCRNN model [3] using a distance based adjacency matrix.
  4. 4.
    GWNET: The GWNET model [4] using a self-adaptive adjacency matrix, where the graph connects regions within 4 kilometres.
  5. 5.
    TAU: The TAU model [5].
  6. 6.
    GATAU: The input convolutional layer is composed of two sequential 2D convolution layers with a hidden size 16. The TAU units have the same model parameters as in the original work [5]. Our output layer is composed of two 2D convolution layers at output with hidden size 16, and a kernel size (1,1), one final linear layer to match dimensions to the output dimension. For the Graph Attention layer, the graph connects regions within 4 kilometres.
@@ V-D Baselines
  2. 2.
    FedAvg: FedAvg [7] uses a weighted averaging mechanism, based on relative client dataset sizes, to update the global model based on individual client contributions.
  3. 3.
    FedOpt: FedOpt [21] builds on FedAvg by implementing server-side adaptive optimisers such as Adam, as well as server-side momentum. For the server, we use a learning rate of 0.01, a momentum parameter of 0.9, and a second moment parameter of 0.99.
  4. 4.
    FedProx: FedProx [22] utilises a penalty term, $\mu$, to inhibit deviations of clients from the global model. We use a proximal term $\mu$ of 0.001.
  5. 5.
@@
The code for this project is made available at https://github.com/FerOroC/FedTPS_Traffic_Prediction.
VI Results
VI-A Traffic Flow Data Generation
Figure 6 depicts heatmaps of traffic flows generated by the TFDiff compared with real traffic flow data, as well as the total region traffic flows over time intervals on a Monday.
Figure 6: Heatmap visualisation of real and generated synthetic traffic flow data for the city of Chengdu from the federated TFDiff model, as well as the total traffic volumes for real and synthetic traffic flow data over all time intervals on a Monday.
Table III details the performance of various generative models based on three accuracy metrics which compare the similarity between the generated data, and the real data. The generative ability of the TFDiff diffusion model to generate traffic data with the same spatio-temporal characteristics as the real data is illustrated by the much lower MAD, and the very high structure similarity and temporal correlation metrics. Although the CGAN and CVAE models achieved high accuracy scores on the SSI, the high MAD and low TC scores demonstrate a high discrepancy between the real and synthetic data. This is due to the fact that cVAE and cGAN primarily capture spatial distribution, which may result in insufficient representation of the temporal dependencies inherent in traffic flow data.
TABLE III: Comparison of generated regional inflow data against real regional inflow data for various centralized generative models.
City Model MAD SSI TC
Chengdu CGAN 0.53 0.86 0.33
CVAE 0.55 0.93 0.53
TFDiff 0.14 0.97 0.92
Xi’an CGAN 1.02 0.78 0.22
CVAE 1.40 0.88 0.29
TFDiff 0.26 0.93 0.91
For the TFDiff model trained under federated settings with a varying number of clients, Table IV shows the similarity between generated samples and real traffic data based on four metrics, averaged over the regional inflow data sequences. Some deviation is expected due to the generational diversity of the DiffTraj model, yet the high SSI and TC metrics demonstrate that the spatio-temporal characteristics of the original data were modelled accurately by the trajectory generation model. In general, increasing the number of clients in the FL setting results in a larger deviation from the real traffic data for the samples generated, as evidenced by the decreasing values of SSI and Temporal Correlation, as well as increasing nMAE and MAPE metrics.
TABLE IV: Comparison of generated regional inflow data against real regional inflow data for the TFDiff model trained through FL, with varying client cases.
City Clients nMAE MAPE SSI TC
Chengdu 2 0.13 16.67 0.97 0.91
4 0.18 21.21 0.93 0.86
6 0.20 22.26 0.91 0.82
8 0.23 23.77 0.88 0.79
Xi’an 2 0.31 31.25 0.94 0.88
4 0.33 29.04 0.91 0.85
6 0.38 29.90 0.88 0.81
8 0.43 31.47 0.86 0.78
TABLE V: Comparison of various Spatio-Temporal model performances for prediction of traffic at 6 time-steps in future. Assessed using nMAE and MAPE metrics, under varying client numbers and FL methods. (All numeric values are shown to 4 significant figures.)